OpenStudy (anonymous):

A new movie open the other day. So far 500,000 people have seen. The producers of the movie need to know if the people liked it. A random sample of 8000 people were asked as they were leaving the theater if they like the movie. Of those interviewed 4200 enjoyed the movie. Predict the total number of people who have enjoyed the movie.

OpenStudy (anonymous):

\[\frac{ 500,000 }{ 8000 }=\frac{ x }{ 4200 }\]\[500,000\times4200=2,100,000,000\]\[8000\times x=8000x\]\[8000x \div 8000=x\]\[2,100,000,000\div8000=262,500\]So, 262,500 people could have liked the movie.

OpenStudy (anonymous):

Thank so much

OpenStudy (anonymous):

Your welcome.